OpenStudy (anonymous):

f(x) = 4x-8 g(x) = x+4 / 8 find f(g(x)) and g(f(x)) and if they are inverses of each other

OpenStudy (anonymous):

my answers are \[x-12\div2 \] and \[x-1\div2\] and they are not inverses of each other. correct?

OpenStudy (anonymous):

answers are correct but they are not inverses of each other

OpenStudy (anonymous):

lets find inverse of y=(x-12)/2

OpenStudy (anonymous):

swap the x and y and solve for y what will u get?

OpenStudy (anonymous):

i dont need to find the inverses, just needed to know that they werent. but thank you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

heres another one though, one sec

OpenStudy (anonymous):

f(g(x))=(x-12)/2 g(f(x))=(x-1)/2 and \[f^(-1)(x)=(x+8)/4\] \[g ^{-1}(x)=8x-4\] so these functions not inverses each others

OpenStudy (anonymous):

 starts at 5, highest point is 3 cross y acis at 2 and crosses x axis again at -2